OpenStudy (anonymous):

help? i dont know what thje question is asking for.. ? :/ Evaluate P(6, 2). a 12 b30 c360

OpenStudy (anonymous):

mabe the product of 6 and 2 which would be 12...but only maybe

OpenStudy (anonymous):

It seems a questions on combinatorics and the answer then is 30. 6*5 = 30

OpenStudy (anonymous):

other examples P(6,3) = 6*5*4 P(9,4)=9*8*7*6 P(100,2) = 100*99